OpenStudy (anonymous):

solve 2/z = 8/z - 1/2

OpenStudy (anonymous):

6/z = 1/2z = 12

OpenStudy (anonymous):

sorry z = 12

OpenStudy (anonymous):

do u follow it ok?

OpenStudy (anonymous):

yeah, i had gotten as far as your first step but was unsure how to procede, thank you!